# FoC: Figure out the Cryptographic Functions in Stripped Binaries with LLMs

### v2

@@ 9.1. Binary Analysis
The integration of these deep learning methods expands the potential of binary analysis, in order to cope with increasingly complex software systems. In this work, we mainly focus on two aspects: binary code summarization and binary code similarity detection.
@@ 9.2. LLM for SE
The emergence and rapid development of Large Language Models (LLMs) have triggered disruptive changes across many fields, including software engineering (SE) (hou2023large, ; zhang2023survey, ). An increasing number of research works are exploring the application of LLMs in various software engineering tasks, with their impact spanning across the the entire software development life cycle, from requirements engineering (sridhara2023chatgpt, ), software design (mandal2023large, ) and development (dong2023self, ), software quality assurance (tang2023csgvd, ), to software maintenance and management (alhamed2022evaluation, ).

### v3

@@ 9.1. Binary Analysis
With the rapid development of Large Language Model (LLM) technology, its exceptional performance in natural language processing and program understanding has prompted researchers to explore its application to binary code analysis related tasks. For instance, DeGPT (hu2024degpt, ) optimizes the readability and accuracy of decompilation results by using an LLM, like ChatGPT, to post-process the output of traditional decompilers. Nova (jiang2024nova, ), built upon the DeepSeek-Coder model (guo2024deepseek, ), introduces a hierarchical attention mechanism and contrastive learning techniques to better capture the high-level semantics of binary code, significantly improving accuracy and interpretability in the analysis of complex binary programs. Similarly, LLM4Decompile (tan2024llm4decompile, ), which is also based on DeepSeek-Coder, is fine-tuned specifically for binary code decompilation tasks. Meanwhile, the Machine Language Model (MLM) (MLM, ), as a closed-source LLM product specifically for machine language, has demonstrated great potential in the structural analysis, semantic analysis, and security analysis of binary programs, addressing challenges such as missing binary information and difficulties in semantic comprehension.
The integration of these deep learning methods and large language models expand the potential of binary analysis, in order to cope with increasingly complex software systems. In this work, we mainly focus on two aspects: binary code summarization and binary code similarity detection.
@@ 9.2. LLM for SE
The emergence and rapid development of Large Language Models (LLMs) have triggered disruptive changes across many fields, including software engineering (SE) (hou2023large, ; zhang2023survey, ). An increasing number of research works are exploring the application of LLMs in various software engineering tasks, with their impact spanning across the the entire software development life cycle, from requirements engineering (sridhara2023chatgpt, ), software design (mandal2023large, ) and development (dong2023self, ), software quality assurance (tang2023csgvd, ), to software maintenance and management (alhamed2022evaluation, ; zhang2024context, ).